# How to Purchase Labels? A Cost-effective Approach Using Active Learning Markets

### v1

@@ Appendix A Monte Carlo results
This appendix reports additional Monte Carlo simulation results that complement the main estimation results presented in Sections 4.1.5 (Table 4) and 4.2.5 (Table 5). For each configuration, we vary key model parameters—namely willingness-to-pay (WTP) $\phi$, willingness-to-sell (WTS) scaling coefficient $\eta$, and budget limit BB—and report the mean and interquartile range of purchased data points across 50 independent runs. These results illustrate the robustness of the proposed pricing schemes and their sensitivity to parameter variations.
@@
BC SC
Parameter / Value VBAL QBCAL RSC VBAL QBCAL RSC
@@
10 7.50 [7.00–8.00] 7.84 [7.00–8.00] 3.54 [1.00–4.50] 7.48 [7.00–8.00] 7.10 [6.00–8.00] 13.56 [14.00–17.00]
15 7.74 [7.00–8.00] 8.10 [7.00–8.75] 5.76 [1.00–14.00] 7.66 [7.00–8.00] 7.66 [7.00–8.00] 15.32 [15.00–20.00]
20 7.74 [7.00–8.00] 8.10 [7.00–8.75] 8.14 [1.00–15.75] 7.72 [7.00–8.00] 7.90 [7.00–8.75] 15.38 [15.00–20.00]
25 7.74 [7.00–8.00] 8.10 [7.00–8.75] 12.94 [9.50–17.75] 7.72 [7.00–8.00] 7.98 [7.00–8.75] 15.38 [15.00–20.00]
30 7.74 [7.00–8.00] 8.10 [7.00–8.75] 15.08 [14.25–19.75] 7.74 [7.00–8.00] 8.00 [7.00–8.75] 15.38 [15.00–20.00]
@@
BC SC
Parameter / Value VBAL QBCAL RSC VBAL QBCAL RSC
@@
800 7.42 [5.00–9.75] 8.02 [6.00–9.75] 53.94 [20.25–68.00] 7.76 [6.00–9.00] 6.86 [5.00–9.00] 14.90 [13.00–16.75]
1000 9.20 [6.00–12.75] 9.56 [6.00–12.00] 70.90 [25.25–125.50] 8.46 [7.00–10.00] 7.62 [5.25–10.00] 18.72 [16.00–20.00]
1200 10.46 [7.00–13.75] 10.98 [7.00–14.00] 88.88 [35.00–141.50] 9.00 [7.00–11.00] 8.20 [5.25–11.00] 22.18 [20.00–24.00]
1400 11.44 [8.25–14.00] 11.90 [8.00–14.75] 106.36 [46.50–154.00] 9.50 [7.00–11.75] 8.68 [6.00–11.00] 25.36 [23.00–28.00]
1600 12.12 [9.25–15.00] 12.56 [9.00–16.00] 116.78 [57.00–159.75] 9.82 [7.25–12.00] 9.08 [7.00–11.75] 28.80 [26.00–31.75]

### v2

@@ Appendix A Monte Carlo results
This appendix reports additional Monte Carlo simulation results that complement the main estimation results presented in Sections 4.1.5 (Table 4) and 4.2.5 (Table 5). For each configuration, we vary key model parameters—namely willingness-to-pay (WTP) $\phi$, willingness-to-sell (WTS) scaling coefficient $\eta$, and budget limit BB—and report the mean and interquartile range of purchased data points across 50 independent runs.
@@
BC
Parameter / Value VBAL QBCAL RSC GreedyKnapsack
@@
10 8.14 [8.00–9.00] 8.42 [8.00–9.00] 17.88 [16.00–20.00] 28.84 [25.25–33.00]
15 8.36 [8.00–9.00] 8.66 [8.00–9.00] 18.34 [16.00–20.75] 29.74 [27.00–33.00]
20 8.36 [8.00–9.00] 8.66 [8.00–9.00] 18.34 [16.00–20.75] 29.74 [27.00–33.00]
25 8.36 [8.00–9.00] 8.66 [8.00–9.00] 18.34 [16.00–20.75] 29.74 [27.00–33.00]
30 8.36 [8.00–9.00] 8.66 [8.00–9.00] 18.34 [16.00–20.75] 29.74 [27.00–33.00]
SC
Parameter / Value VBAL QBCAL RSC GreedyKnapsack
WTP $\phi$ [£/TWD2]
1000 8.06 [7.25–9.00] 7.44 [6.25–8.75] 18.30 [16.00–20.75] 29.74 [27.00–33.00]
1050 8.14 [8.00–9.00] 7.74 [7.00–9.00] 18.30 [16.00–20.75] 29.74 [27.00–33.00]
1100 8.16 [8.00–9.00] 7.94 [7.00–9.00] 18.30 [16.00–20.75] 29.74 [27.00–33.00]
1150 8.26 [8.00–9.00] 8.08 [7.00–9.00] 18.30 [16.00–20.75] 29.74 [27.00–33.00]
1200 8.34 [8.00–9.00] 8.16 [7.00–9.00] 18.30 [16.00–20.75] 29.74 [27.00–33.00]
WTS scaling coefficient $\eta$
0.5 8.34 [8.00–9.00] 8.16 [7.00–9.00] 18.30 [16.00–20.75] 29.74 [27.00–33.00]
1 6.40 [4.00–8.00] 4.68 [3.00–7.00] 14.36 [13.00–16.00] 27.48 [26.00–29.00]
1.5 4.30 [2.00–6.75] 2.80 [1.00–4.75] 10.00 [9.00–11.00] 24.06 [23.00–26.00]
2 3.10 [2.00–4.00] 2.10 [1.00–3.00] 7.74 [7.00–9.00] 21.70 [21.00–23.00]
2.5 2.48 [2.00–3.00] 1.60 [1.00–3.00] 6.40 [6.00–7.00] 18.88 [18.00–20.00]
Budget limit $B$ [£]
10 8.28 [8.00–9.00] 7.72 [7.00–9.00] 16.44 [14.25–18.00] 28.22 [27.00–31.00]
15 8.34 [8.00–9.00] 8.16 [7.00–9.00] 18.30 [16.00–20.75] 29.74 [27.00–33.00]
20 8.36 [8.00–9.00] 8.32 [7.00–9.00] 18.34 [16.00–20.75] 29.74 [27.00–33.00]
25 8.36 [8.00–9.00] 8.38 [7.00–9.00] 18.34 [16.00–20.75] 29.74 [27.00–33.00]
30 8.36 [8.00–9.00] 8.56 [7.25–9.00] 18.34 [16.00–20.75] 29.74 [27.00–33.00]
@@
BC
Parameter / Value VBAL QBCAL RSC
@@
800 7.42 [5.00–9.75] 8.02 [6.00–9.75] 53.94 [20.25–68.00]
1000 9.20 [6.00–12.75] 9.56 [6.00–12.00] 70.90 [25.25–125.50]
1200 10.46 [7.00–13.75] 10.98 [7.00–14.00] 88.88 [35.00–141.50]
1400 11.44 [8.25–14.00] 11.90 [8.00–14.75] 106.36 [46.50–154.00]
1600 12.12 [9.25–15.00] 12.56 [9.00–16.00] 116.78 [57.00–159.75]
SC
Parameter / Value VBAL QBCAL RSC
WTP $\phi$ [£/TWD2]
30 6.62 [5.00–8.00] 5.80 [4.00–8.00] 22.18 [20.00–24.00]
40 7.96 [6.00–10.00] 7.02 [5.00–9.00] 22.18 [20.00–24.00]
50 9.00 [7.00–11.00] 8.20 [5.25–11.00] 22.18 [20.00–24.00]
60 9.86 [8.00–12.00] 8.84 [6.00–12.00] 22.18 [20.00–24.00]
70 10.44 [8.00–12.75] 9.40 [7.00–12.00] 22.18 [20.00–24.00]
WTS scaling coefficient $\eta$
1 9.00 [7.00–11.00] 8.20 [5.25–11.00] 22.18 [20.00–24.00]
2 4.56 [3.00–6.00] 3.94 [3.00–5.00] 11.02 [10.00–12.00]
3 2.78 [1.25–4.00] 2.44 [1.00–3.00] 7.52 [6.00–9.00]
4 1.76 [1.00–3.00] 1.60 [1.00–2.00] 5.22 [4.00–6.00]
5 1.20 [0.00–2.00] 1.14 [0.00–2.00] 4.18 [3.00–5.00]
Budget limit $B$ [£]
800 7.76 [6.00–9.00] 6.86 [5.00–9.00] 14.90 [13.00–16.75]
1000 8.46 [7.00–10.00] 7.62 [5.25–10.00] 18.72 [16.00–20.00]
1200 9.00 [7.00–11.00] 8.20 [5.25–11.00] 22.18 [20.00–24.00]
1400 9.50 [7.00–11.75] 8.68 [6.00–11.00] 25.36 [23.00–28.00]
1600 9.82 [7.25–12.00] 9.08 [7.00–11.75] 28.80 [26.00–31.75]